OpenStudy (anonymous):

how would i find x-y when x=2.5 and y=0.06

OpenStudy (anonymous):

you have your values for what x and way are plug them into the fourmula and do the math.

OpenStudy (anonymous):

x-y = 2.5 - 0.06 =2.44